OpenStudy (anonymous):

convert 5pi/7 into degrees

OpenStudy (anonymous):

multiply that by 180 degrees, then divide by pi

OpenStudy (anonymous):

\[degrees=radians*\frac{ 180 }{ \pi }\]

OpenStudy (anonymous):

\[\frac{ 5\pi }{ 7 } \times \frac{ 180 }{ \pi }\]